# The Dirichlet heat trace for domains with curved corners

### v1

@@ Abstract.
We study the short-time asymptotics of the Dirichlet heat trace on planar curvilinear polygons. For such domains we show that the coefficient of $t^{1/2}$ in the expansion splits into a boundary integral of $\kappa^{2}$ and a sum of local corner contributions, one for each vertex. Each curved corner contribution depends only on the interior angle $\alpha$ and on the limiting curvatures $\kappa_{\pm}$ on the adjacent sides. Using a conformal model and a parametrix construction on the sector heat space, we express this contribution in the form $c_{1/2}(\alpha)\,r_{0}(\alpha,\kappa_{+},\kappa_{-})$, where $c_{1/2}(\alpha)$ is given by a Hadamard finite part of an explicit trace over the exact sector. For right-angled corners we compute $c_{1/2}(\pi/2)=1/(16\sqrt{\pi})$ and obtain a closed formula for the $t^{1/2}$ coefficient. As an application we extend a previous result in the literature by showing that any admissible curvilinear polygon that is Dirichlet isospectral to a polygon must itself be a polygon with straight sides.
@@ 1. Introduction
A classical problem in spectral geometry is to determine what geometric information is encoded in the spectrum of the Laplacian on a Riemannian manifold [kac1966can]. This is frequently approached using heat trace methods. Specifically, one examines the short-time asymptotic expansion of the trace of the heat kernel. This trace is spectrally determined, and therefore the coefficients in the asymptotic expansion are spectral invariants – geometric quantities determined by the spectrum. In 1967, McKean and Singer computed the first three heat invariants for a smoothly bounded domain in $\mathbb{R}^{2}$ [mckean1967curvature]. Using these, they showed that the area, perimeter, and Euler characteristic of such a domain are all spectral invariants. It is natural to ask whether one can deduce similar information for a less smooth domain, and indeed this has been studied in depth [Cheeger1983, vanDenBergSrisat88, NRS].
@@
It is well known that the trace of the Dirichlet heat kernel on $\Omega$, $H^{D}_{\Omega}(t)$, has an asymptotic expansion as $t\to 0$, beginning with a $t^{-1}$ term. The $t^{-1}$, $t^{-1/2}$, and $t^{0}$ terms of this expansion have been computed, first in the exact polygonal setting in unpublished work of Ray, as cited by [Cheeger1983] and [vdBS], and then extended to the general curvilinear setting [NRS]. Our interest here is in the next term, which we expect at order $t^{1/2}$. In [EGS_2017], based on results in the literature such as [branson1990asymptotics], the authors give the coefficient of this $t^{1/2}$ term in the “straight corners” case, where a neighborhood of each corner $P_{j}$ is isometric to an exact sector. Without this assumption, the curvature of the sides may interact with the angle at the corner. As we show, this interaction produces a term at order $t^{1/2}$.
@@
(2) $r_{0}=r_{0}(\alpha,\kappa_{+},\kappa_{-})=\frac{1}{2}\csc\alpha\sqrt{\kappa_{+}^{2}+\kappa_{-}^{2}-2\kappa_{+}\kappa_{-}\cos\alpha}.$
@@
$\quad a\left(\frac{R}{\sqrt{1-\sigma}},\theta,\frac{\sqrt{\sigma}\,\rho}{\sqrt{1-\sigma}},\phi\right)\,b\left(\rho,\phi,\frac{R}{\sqrt{\sigma}},\theta\right)\,\,\rho\,d\rho\,d\phi\,\,R\,dR\,d\theta.$
@@
(5) $c_{1/2}(\frac{\pi}{2})=\frac{1}{16\sqrt{\pi}}.$
@@ 1.1. Inverse spectral applications
In [EGS_2017], the authors prove that any curvilinear polygon with straight corners is in fact isospectral to a polygon. They require straight corners because of a lack of a formula in the curved corner case. We can strengthen their result.
@@
A curvilinear polygon $\Omega$ is admissible if, for each $j$, either $c_{1/2}(\alpha_{j})>0$ or $\Omega$ is straight near $P_{j}$.
@@
(6) $\frac{1}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds+\sum_{j=1}^{n}c_{1/2}(\alpha_{j})\sqrt{\kappa_{+,j}^{2}+\kappa_{-,j}^{2}}.$
Since this coefficient is non-negative and only zero for an exact polygon, $\Omega$ is an exact polygon. ∎
@@ 1.2. Plan of the paper
In section 2 we prove Theorem 1. This is done by examining the geometric microlocal description of the heat kernel given in [NRS]. We show that there is a well-defined corner contribution $\mathcal{C}_{1/2}\sqrt{t}$, and then prove that any two corners with the same angle $\alpha$ and limit curvatures $\kappa_{\pm}$ yield the same contribution, regardless of the rest of $\Omega$. To take advantage of this uniformity, in section 3 we introduce a simple conformal model for a curvilinear corner and compute the contribution $\mathcal{C}_{1/2}\sqrt{t}$ in that model. This computation proceeds via a parametrix construction as in [NRS] and allows us to obtain both (3) and (4). Finally, in section 4, we prove (5) by computing a sufficient number of terms in the Dirichlet heat trace expansion for a semicircle and comparing with (1).
@@ Acknowledgements
D. S. is grateful for support from the AMS-Simons research enhancement grant 501949-9208. S. L. is grateful for support from NSF grant DMS-2346799.
@@ 2. The corner contribution
Our approach to understanding the heat kernel on $\Omega$ is based on the methods of geometric microlocal analysis [MelroseAPS, NRS]. The strategy is to first resolve the singularities of the heat kernel at $t=0$ by constructing a “double heat space,” $\Omega_{h}^{2}$, via a sequence of geometric blow-ups. The heat kernel lifts under these blow-ups to a well-behaved (“polyhomogeneous conormal”) function on the manifold with corners $\Omega_{h}^{2}$ [NRS]. Roughly speaking, polyhomogeneous conormal functions are those whose asymptotic expansions at the boundary hypersurfaces are compatible and jointly polyhomogeneous at higher-codimension corners; the precise definition may be found in [MelroseAPS].
@@
The method of proof in [NRS], which is important for our work here, is to use geometric microlocal analysis to perform a parametrix construction. Specifically, one constructs a parametrix $K(t,z,z^{\prime})$, an approximate solution to the heat equation, by specifying its leading order behavior at each of the three boundary hypersurfaces td, sf, and ff. This should be done in such a way so that when the heat operator is applied to the parametrix $K(t,z,z^{\prime})$, the result $R(t,z,z^{\prime})$ is lower order at td, sf, and ff. In order to arrange this, we choose leading order behavior that solves the relevant model problems for the heat equation at each boundary hypersurface. We also must ensure that the specified behaviors at td, sf, and ff are consistent with each other at the intersections of the boundary hypersurfaces, so that there actually exists a polyhomogeneous conormal function on $\Omega_{h}^{2}$ with all of the correct behavior. Once we have a good enough parametrix, we can recover the true heat kernel via the Volterra series
@@
To analyze this Volterra series, we use a composition formula for polyhomogeneous conormal functions on $\Omega_{h}^{2}$, which in our setting is [NRS, Theorem 3.16].
@@
The model cone heat kernel has order 2-2 at td, sf, and ff. We write $\operatorname{ord}_{\mathrm{ff}}(\cdot)$ for this $T$-order. The key result of [NRS, Theorem 3.16] for our setting is that if $B$ vanishes to infinite order at td, then $\operatorname{ord}_{\mathrm{ff}}(A\star B)=\operatorname{ord}_{\mathrm{ff}}(A)+\operatorname{ord}_{\mathrm{ff}}(B)+4$.
As a first use of these techniques, we show that the corner contribution $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})$ indeed depends only on $\alpha$ and $\kappa_{\pm}$. Suppose that there exist two domains $\Omega$ and $\Omega^{\prime}$, each with one corner, with the same $\alpha$ and the same $\kappa_{\pm}$. By [NRS], the heat kernels $H^{D}_{\Omega}$ and $H^{D}_{\Omega^{\prime}}$ are polyhomogeneous conormal on $\Omega_{h}^{2}$ and $(\Omega^{\prime})_{h}^{2}$ respectively. Since $\alpha$ is the same, the faces ff of these respective heat spaces are geometrically identical. We now state a lemma which informally says that: under these conditions, $H^{D}_{\Omega}-H^{D}_{\Omega^{\prime}}$ has leading order at worst $t^{0}$ at ff.
@@
  • At sf, we specify that $K$ should have the same expansion as the Dirichlet heat kernel on a smooth domain with boundary whose boundary coincides with the appropriate smooth part of $\partial\Omega$. Such a heat kernel is known to have a polyhomogeneous conormal expansion by [Grieser], and that expansion is consistent with the Euclidean heat kernel at td. The first term, at order 2-2, is universal. The second term, at order 1-1, follows from the boundary calculus construction in [Grieser]; by a standard scaling argument, this coefficient is a universal multiple of the curvature $\kappa(z^{\prime})$.
@@
This consistency means that there exists a parametrix $K(t,z,z^{\prime})$ with the specified expansions at td, sf, and ff. This parametrix solves the model problems to infinite order at td and sf and to two orders at ff. So $R(t,z,z^{\prime})=(\partial_{t}-\Delta_{\Omega})K(t,z,z^{\prime})$ has leading order 2-2, as opposed to 4-4, at ff, and leading order $\infty$ at td and sf. Using the composition formula of [NRS], we conclude that for each $n\geq 1$, $K\star(R^{\star n})$ has leading order at worst $t^{0}$ at ff.
@@
has order at worst 0 at ff, and since the same is true for $H^{D}_{\Omega^{\prime}}-K$, so $H^{D}_{\Omega}-H^{D}_{\Omega^{\prime}}=O(T^{0})$ at ff. The Lemma is proven. ∎
@@
As in [NRS], the heat trace is obtained by restricting $H^{D}_{\Omega}$ to the spatial diagonal and then integrating over $z\in\Omega$. The technical tool used is Melrose’s pushforward theorem. We will not repeat all the details here, but the upshot is that each face td, sf, and ff gives a separate contribution to the expansion (1). A priori, the expansions can interact to give logarithmic terms, but for the same reasons as in [NRS, p. 49], keeping in mind that we now have two orders at ff rather than one, there are no logarithmic terms until at least $O(T^{2}\log T)=O(t\log t)$ (and potentially much later, if at all). The contribution from td is

### v2

# The heat trace for domains with curved corners

@@ Abstract.
We derive the short-time expansion of the heat trace on planar curvilinear polygons through order $t^{1/2}$, with both Dirichlet and Neumann boundary conditions. For such domains we show that the coefficient of $t^{1/2}$ in the expansion splits into a boundary integral of $\kappa^{2}$ and a sum of local corner contributions, one for each vertex. Each curved corner contribution depends only on the interior angle $\alpha$ and on the limiting curvatures $\kappa_{\pm}$ on the adjacent sides. For Dirichlet boundary conditions, using a conformal model and a parametrix construction on the sector heat space, we express this contribution in the form $c_{1/2}(\alpha)\,r_{0}(\alpha,\kappa_{+},\kappa_{-})$, where $c_{1/2}(\alpha)$ is given by the Hadamard finite part of an explicit trace over the exact sector. For right angles we compute $c_{1/2}(\pi/2)=1/(16\sqrt{\pi})$, which yields a closed formula for the $\sqrt{t}$ coefficient for domains with right-angled corners. Combined with previous work on the coefficients of $t^{-1}$, $t^{-1/2}$ and $t^{0}$ for such domains, this determines the remaining local heat trace coefficient for planar curvilinear polygons up to order $t^{1/2}$. As an inverse spectral application we extend a result in the literature by showing that any admissible curvilinear polygon that is Dirichlet isospectral to a polygon must itself be a polygon with straight sides.
@@ 1. Introduction
A classical problem in spectral geometry is to determine what geometric information is encoded in the spectrum of the Laplacian on a Riemannian manifold [8]. This is frequently approached using heat trace methods. Specifically, one examines the short-time asymptotic expansion of the trace of the heat kernel. This trace is spectrally determined, and therefore the coefficients in the asymptotic expansion are spectral invariants – geometric quantities determined by the spectrum. In 1967, McKean and Singer computed the first three heat invariants for a smoothly bounded domain in $\mathbb{R}^{2}$ [7]. Using these, they showed that the area, perimeter, and Euler characteristic of such a domain are all spectral invariants. It is natural to ask whether one can deduce similar information for a less smooth domain, and indeed this has been studied in depth [2, 11, 10].
@@
It is well known that the trace of the heat kernel on $\Omega$ with either Dirichlet or Neumann boundary conditions, $H^{D}_{\Omega}(t)$ or $H^{N}_{\Omega}(t)$ respectively, has an asymptotic expansion as $t\to 0$, beginning with a $t^{-1}$ term. The $t^{-1}$, $t^{-1/2}$, and $t^{0}$ terms of this expansion have been computed, first in the exact polygonal setting in unpublished work of Ray, as cited in [2, 12], and then extended to the general curvilinear setting in [10]. The next local heat invariant is the coefficient of $t^{1/2}$; it is the first term beyond the regime treated in [2, 11, 10] and the first coefficient that couples boundary curvature to the corner geometry. In [4], using results such as [1], this $t^{1/2}$ coefficient is computed in the “straight corners” case, where a neighborhood of each corner $P_{j}$ is isometric to an exact sector. Without this assumption, the curvature of the sides may interact with the angle at the corner. As we show, this interaction produces a term at order $t^{1/2}$.
Our first main result identifies this $t^{1/2}$ coefficient, showing that it splits into a boundary integral of $\kappa^{2}$ and a sum of local corner terms, one for each vertex. It thus completes the description of the heat trace asymptotics up to order $t^{1/2}$ for planar curvilinear polygons. We also obtain this result with Neumann boundary conditions, which we state in Proposition 1.
@@
(2) $r_{0}=r_{0}(\alpha,\kappa_{+},\kappa_{-})=\frac{1}{2}|\csc\alpha|\sqrt{\kappa_{+}^{2}+\kappa_{-}^{2}+2\kappa_{+}\kappa_{-}\cos\alpha}.$
@@
$\quad a\left(\frac{R}{\sqrt{1-\sigma}},\theta,\frac{\sqrt{\sigma}\,\rho}{\sqrt{1-\sigma}},\phi\right)\,b\left(\rho,\phi,\frac{R}{\sqrt{\sigma}},\theta\right)\,\,\rho\,d\rho\,d\phi\,\,R\,dR\,d\theta\,d\sigma.$
@@
(5) $c_{1/2}\left(\frac{\pi}{2}\right)=\frac{1}{16\sqrt{\pi}}.$
The factorization (3) shows that, away from $\alpha=\pi$, the corner contribution at order $t^{1/2}$ splits cleanly into an angular factor $c_{1/2}(\alpha)$ and a curvature factor $r_{0}(\alpha,\kappa_{+},\kappa_{-})$, so that all dependence on the limiting curvatures is encoded in a single scalar quantity. This theorem is not just an integral identity, but a structural description of how corner curvature enters the expansion.
@@
  • In (1), the corner contributions to the coefficients of $t^{-1}$ and $t^{0}$ depend only on the interior angles $\alpha_{j}$; the one-sided curvatures $\kappa_{j,\pm}$ do not appear. Theorem 2 shows that the first place where the corner contribution becomes sensitive to $(\kappa_{j,+},\kappa_{j,-})$ is at order $t^{1/2}$.
1.0.1. The Neumann case
Theorem 1 may be generalized, via identical arguments, to the case of the Neumann heat trace $H^{N}_{\Omega}(t)$. The first four terms of the Neumann heat trace expansion in the straight-corners setting are given in [4]. We obtain
Proposition 1.
Under the same assumptions as in the Dirichlet case,
(6) $H^{N}_{\Omega}(t)=\frac{|\Omega|}{4\pi t}+\frac{|\partial\Omega|}{8\sqrt{\pi t}}+\frac{1}{12\pi}\Big(\int_{\partial\Omega}\kappa\,ds+\sum_{j=1}^{n}\frac{\pi^{2}-\alpha_{j}^{2}}{2\alpha_{j}}\Big)\\ +\sqrt{t}\Big(\frac{5}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds+\sum_{j=1}^{n}\mathcal{C}^{N}_{1/2}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})\Big)+O(t\log t).$
It is likely that a theorem of the form Theorem 2 holds in the Neumann case as well. However, there are some complications in the arguments in section 3 that make it more difficult to prove.
@@ 1.1. Inverse spectral applications
In [4], Enciso and Gómez-Serrano prove that any curvilinear polygon with straight corners is in fact isospectral to a polygon. They require straight corners because of a lack of a formula in the curved corner case. Using the description of the $t^{1/2}$ coefficient above, we can strengthen their result.
@@
A curvilinear polygon $\Omega$ is admissible if, for each corner $P_{j}$ with opening angle $\alpha_{j}$, either
$c_{1/2}(\alpha_{j})>0\quad\text{or}\quad\kappa_{j,+}=\kappa_{j,-}=0.$
@@
(7) $\frac{1}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds+\sum_{j=1}^{n}c_{1/2}(\alpha_{j})r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})$
with $r_{0}$ defined in (2). Note that $\int_{\partial\Omega}\kappa^{2}\,ds\geq 0$. Moreover $r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})\geq 0$ by definition, and (for $\alpha_{j}\neq\pi$) one has $r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})=0$ if and only if $\kappa_{j,+}=\kappa_{j,-}=0$. By admissibility, whenever $r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})>0$ we necessarily have $c_{1/2}(\alpha_{j})>0$. Hence every term in (7) is non-negative. Since the $t^{1/2}$ coefficient vanishes, each term must vanish: $\int_{\partial\Omega}\kappa^{2}\,ds=0$ and $r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})=0$ for all $j$. Therefore $\kappa\equiv 0$ on each smooth edge and $\kappa_{j,+}=\kappa_{j,-}=0$ at every corner, so $\partial\Omega$ is piecewise linear and $\Omega$ is a polygon. ∎
Remark.
The admissibility hypothesis is used only to ensure that whenever a corner has nonzero curvature on at least one side, the corresponding coefficient satisfies $c_{1/2}(\alpha_{j})>0$. The alternative condition $\kappa_{j,+}=\kappa_{j,-}=0$ requires only vanishing one-sided curvature limits; it does not assume that the boundary is straight in a full neighborhood of $P_{j}$.
@@
Together with the known coefficients of order $t^{-1}$, $t^{-1/2}$ and $t^{0}$, Theorems 1 and 2 show that, for planar curvilinear polygons, all local heat invariants up to order $t^{1/2}$ can be expressed in explicit geometric terms and used in inverse spectral arguments of this type.
Remark (A sign condition that would enlarge the admissible class).
In our argument, admissibility is used only through the implication
$(\kappa_{j,+},\kappa_{j,-})\neq(0,0)\quad\Longrightarrow\quad c_{1/2}(\alpha_{j})>0$
at each corner $P_{j}$. Any structural information on the sign of $\alpha\mapsto c_{1/2}(\alpha)$ would therefore allow one to weaken the definition of an admissible curvilinear polygon without changing the proofs.
We conjecture that $c_{1/2}(\alpha)>0$ for all $\alpha\in(0,\pi)$. In that case, every curvilinear polygon with corner angles in $(0,\pi)$ would be admissible in our sense.
1.2. Circular and annular sectors
We record two model families in which the coefficient of $t^{1/2}$ in (1) admits a closed expression. These formulas give concrete examples of the contribution of curved corners to the heat trace at order $t^{1/2}$.
For a curvilinear polygon $\Omega$, write
(8) $H^{D}_{\Omega}(t)=\cdots+a_{1/2}(\Omega)\,\sqrt{t}+O(t\log t),\qquad a_{1/2}(\Omega)=\frac{1}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds+\sum_{j=1}^{n}\mathcal{C}_{1/2}(\alpha_{j},\kappa_{j,+},\kappa_{j,-}).$
In the examples below, every curved corner is a right angle. We therefore use the special value
(9) $c_{1/2}\left(\frac{\pi}{2}\right)=\frac{1}{16\sqrt{\pi}}.$
Note that if $\alpha_{j}=\pi/2$, then $\cos\alpha_{j}=0$ and $\csc\alpha_{j}=1$, and (2) reduces to
(10) $r_{0}\left(\frac{\pi}{2},\kappa_{+},\kappa_{-}\right)=\frac{1}{2}\sqrt{\kappa_{+}^{2}+\kappa_{-}^{2}}.$
Combining (9) and (10) gives
(11) $\mathcal{C}_{1/2}\left(\frac{\pi}{2},\kappa_{+},\kappa_{-}\right)=\frac{1}{32\sqrt{\pi}}\sqrt{\kappa_{+}^{2}+\kappa_{-}^{2}}.$
Circular sectors. Let $R>0$ and $\alpha\in(0,2\pi)$, and set
$S(R,\alpha)=\{(r,\theta)\in\mathbb{R}^{2}:\ 0<r<R,\ 0<\theta<\alpha\}.$
The boundary of $S(R,\alpha)$ consists of two radial segments and one circular arc. There are three corners: the apex with opening angle $\alpha$, and two right-angle corners at the endpoints of the arc.
The basic geometric quantities are
(12) $|S(R,\alpha)|=\frac{\alpha}{2}R^{2},\qquad|\partial S(R,\alpha)|=(\alpha+2)R,\qquad\int_{\partial S(R,\alpha)}\kappa\,ds=\alpha,\qquad\int_{\partial S(R,\alpha)}\kappa^{2}\,ds=\frac{\alpha}{R}.$
The formula (1) then becomes, after a bit of simplification,
(13) $H^{D}_{S(R,\alpha)}(t)=\frac{\alpha R^{2}}{8\pi t}-\frac{(\alpha+2)R}{8\sqrt{\pi t}}+\frac{1}{12\pi}\Big(\frac{3\pi}{2}+\frac{\pi^{2}+\alpha^{2}}{2\alpha}\Big)+a_{1/2}(S(R,\alpha))\sqrt{t}+O(t\log t).$
To find $a_{1/2}$, observe that at the apex, $(\kappa_{+},\kappa_{-})=(0,0)$, hence $\mathcal{C}_{1/2}=0$. At each endpoint of the circular arc, the corner angle is $\pi/2$ and the limiting curvatures are $(\kappa_{+},\kappa_{-})=(1/R,0)$, so (11) gives a contribution $1/(32\sqrt{\pi}R)$. Summing the two endpoints yields
(14) $\sum_{j}\mathcal{C}_{1/2}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})=\frac{1}{16\sqrt{\pi}}\cdot\frac{1}{R}.$
Substituting (12) and (14) into (8) gives
(15) $a_{1/2}\bigl(S(R,\alpha)\bigr)=\frac{1}{256\sqrt{\pi}}\left(\frac{\alpha}{R}\right)+\frac{1}{16\sqrt{\pi}}\left(\frac{1}{R}\right)=\frac{\alpha+16}{256\sqrt{\pi}}\cdot\frac{1}{R},$
which completes our calculation.
Suppose that two circular sectors are Dirichlet isospectral. Then they must be congruent; their congruence can already be deduced from the coefficients of orders $t^{-1}$, $t^{-1/2}$ and $t^{0}$. We give an alternative proof here to show that the coefficients $a_{-1}$ and $a_{1/2}$ are enough by themselves.
Proposition 3 (Spectral determination within the sector family).
[…]
Proof.
Dirichlet isospectrality implies equality of the heat invariants $|\partial\Omega|$ and $a_{1/2}(\Omega)$. Using (12) and (15), we compute
$|\partial S(R,\alpha)|\cdot 256\sqrt{\pi}\,a_{1/2}\bigl(S(R,\alpha)\bigr)=(\alpha+2)R\cdot\frac{\alpha+16}{R}=(\alpha+2)(\alpha+16).$
Hence $\alpha$ is determined as the unique positive solution of
$x^{2}+18x+32=|\partial\Omega|\cdot 256\sqrt{\pi}\,a_{1/2}(\Omega),$
and then $R=|\partial\Omega|/(\alpha+2)$. Uniqueness of $\alpha$ follows since $x^{2}+18x+32$ is strictly increasing on $(0,\infty)$. ∎
Annular sectors. Let
$A(r,R,\alpha)=\{(\rho,\theta)\in\mathbb{R}^{2}:\ r<\rho<R,\ 0<\theta<\alpha\}.$
Then
(16) $|A(r,R,\alpha)|=\frac{\alpha}{2}(R^{2}-r^{2}),\qquad|\partial A(r,R,\alpha)|=\alpha(R+r)+2(R-r).$
[…]
Thus we have
(17) $H^{D}_{A(r,R,\alpha)}(t)=\frac{\alpha(R^{2}-r^{2})}{8\pi t}-\frac{\alpha(R+r)+2(R-r)}{8\sqrt{\pi t}}+\frac{1}{4}+a_{1/2}(A(r,R,\alpha))\sqrt{t}+O(t\log t).$
Using (16),
(18) $a_{1/2}\bigl(A(r,R,\alpha)\bigr)=\frac{\alpha+16}{256\sqrt{\pi}}\left(\frac{1}{R}+\frac{1}{r}\right).$
This additional coefficient may be used to distinguish annuli which have equal heat coefficients up to that point. For example, set
$\Omega_{1}=A\left(1,3,\frac{1}{4}\right),\qquad\Omega_{2}=A\left(2,4,\frac{1}{6}\right).$
A direct check gives
$|\Omega_{1}|=|\Omega_{2}|=1,\qquad|\partial\Omega_{1}|=|\partial\Omega_{2}|=5,$
so $\Omega_{1}$ and $\Omega_{2}$ have the same heat coefficients through $t^{0}$ within this model family. However,
$a_{1/2}(\Omega_{1})=\frac{16+\frac{1}{4}}{256\sqrt{\pi}}\left(1+\frac{1}{3}\right)=\frac{65}{768\sqrt{\pi}},\qquad a_{1/2}(\Omega_{2})=\frac{16+\frac{1}{6}}{256\sqrt{\pi}}\left(\frac{1}{2}+\frac{1}{4}\right)=\frac{97}{2048\sqrt{\pi}},$
so $a_{1/2}(\Omega_{1})\neq a_{1/2}(\Omega_{2})$. Hence $a_{1/2}$ distinguishes noncongruent annular sectors that are indistinguishable by the coefficients through order $t^{0}$.
This said, even after including $a_{1/2}$, the data $(A=|\Omega|,P=|\partial\Omega|,a_{1/2})$ can arise from two distinct triples $(r,R,\alpha)$, hence from two noncongruent annular sectors. To see this, $s=R+r$ and d=Rrd=R-r. We have $A=\frac{\alpha}{2}ds$ and $P=\alpha s+2d$, so eliminating $\alpha$ gives
(19) $P=\frac{2A}{d}+2d.$
For fixed $(A,P)$ with $P>4\sqrt{A}$, the quadratic $2d^{2}-Pd+2A=0$ has two positive roots, so $d$ is not uniquely determined by $(A,P)$. Moreover, from (18) we deduce
(20) $\bigl(256\sqrt{\pi}\,a_{1/2}(\Omega)\bigr)\,(s^{2}-d^{2})=64s+\frac{8A}{d}.$
For each fixed $d>0$, this is a quadratic equation in $s$ with one positive root and one negative root. Thus each admissible choice of $d$ determines a unique $s>d$, and hence a unique triple $(r,R,\alpha)$.
Therefore, even after incorporating the curved-corner invariant in Theorem 2, the data $(A,P,a_{1/2})$ can correspond to two noncongruent annular sectors. To break this remaining ambiguity one would need additional spectral information, for instance higher heat coefficients or another spectral invariant such as the first Dirichlet eigenvalue.
@@ 1.3. Plan of the paper
In section 2 we prove Theorem 1. This is done by examining the geometric microlocal description of the heat kernel given in [10]. We show that there is a well-defined corner contribution $\mathcal{C}_{1/2}\sqrt{t}$, and then prove that any two corners with the same angle $\alpha$ and limit curvatures $\kappa_{\pm}$ yield the same contribution, regardless of the rest of $\Omega$. To take advantage of this uniformity, in section 3 we introduce a simple conformal model for a curvilinear corner and compute the contribution $\mathcal{C}_{1/2}\sqrt{t}$ in that model. This computation proceeds via a parametrix construction as in [10] and allows us to obtain both the factorization (3) and the trace formula (4). Finally, in section 4, we prove (5), i.e. compute $c_{1/2}(\pi/2)$, by computing a sufficient number of terms in the Dirichlet heat trace expansion for a unit half-disk and comparing with (1). Section 4 also contains the same calculation for the Neumann case, identifying $\mathcal{C}^{N}_{1/2}\bigl(\frac{\pi}{2},1,0\bigl)$.
@@ Acknowledgements
D. S. is grateful for support from the AMS-Simons research enhancement grant 501949-9208. S. L. was supported by an Olga Taussky and John Todd Fellowship and by NSF grant DMS-2346799.
@@ 2. The corner contribution
Our approach to understanding the heat kernel on $\Omega$ is based on the methods of geometric microlocal analysis [9, 10]. The strategy is to first resolve the singularities of the heat kernel at $t=0$ by constructing a “double heat space,” $\Omega_{h}^{2}$, via a sequence of geometric blow-ups. The heat kernel lifts under these blow-ups to a well-behaved (“polyhomogeneous conormal”) function on the manifold with corners $\Omega_{h}^{2}$ [10]. Roughly speaking, polyhomogeneous conormal functions are those whose asymptotic expansions at the boundary hypersurfaces are compatible and jointly polyhomogeneous at higher-codimension corners; the precise definition may be found in [9].
@@
The method of proof in [10], which is important for our work here, is to use geometric microlocal analysis to perform a parametrix construction. Specifically, one constructs a parametrix $K(t,z,z^{\prime})$, an approximate solution to the heat equation, by specifying its leading order behavior at each of the three boundary hypersurfaces td, sf, and ff. This should be done in such a way so that when the heat operator is applied to the parametrix $K(t,z,z^{\prime})$, the result $R(t,z,z^{\prime})$ is lower order at td, sf, and ff. In order to arrange this, we choose leading order behavior that solves the relevant model problems for the heat equation at each boundary hypersurface. We also must ensure that the specified behaviors at td, sf, and ff are consistent with each other at the intersections of the boundary hypersurfaces, so that there actually exists a polyhomogeneous conormal function on $\Omega_{h}^{2}$ with all of the correct behavior. Once we have a good enough parametrix, we can recover the true heat kernel via the Volterra series
@@
To analyze this Volterra series, we use a composition formula for polyhomogeneous conormal functions on $\Omega_{h}^{2}$, which in our setting is [10, Theorem 3.16].
@@
The model cone heat kernel has order 2-2 at td, sf, and ff. We write $\operatorname{ord}_{\mathrm{ff}}(\cdot)$ for this $T$-order. The key result of [10, Theorem 3.16] for our setting is that if $B$ vanishes to infinite order at td, then $\operatorname{ord}_{\mathrm{ff}}(A\star B)=\operatorname{ord}_{\mathrm{ff}}(A)+\operatorname{ord}_{\mathrm{ff}}(B)+4$.
As a first use of these techniques, we show that the corner contribution $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})$ indeed depends only on $\alpha$ and $\kappa_{\pm}$. Suppose that there exist two domains $\Omega$ and $\Omega^{\prime}$, each with one corner, with the same $\alpha$ and the same $\kappa_{\pm}$. By [10], the heat kernels $H^{D}_{\Omega}$ and $H^{D}_{\Omega^{\prime}}$ are polyhomogeneous conormal on $\Omega_{h}^{2}$ and $(\Omega^{\prime})_{h}^{2}$ respectively. Since $\alpha$ is the same, the faces ff of these respective heat spaces are geometrically identical. We now state a lemma which informally says that: under these conditions, $H^{D}_{\Omega}-H^{D}_{\Omega^{\prime}}$ has leading order at worst $t^{0}$ at ff.
@@
  • At sf, we specify that $K$ should have the same expansion as the Dirichlet heat kernel on a smooth domain with boundary whose boundary coincides with the appropriate smooth part of $\partial\Omega$. Such a heat kernel is known to have a polyhomogeneous conormal expansion by [6], and that expansion is consistent with the Euclidean heat kernel at td. The first term, at order 2-2, is universal. The second term, at order 1-1, follows from the boundary calculus construction in [6]; by a standard scaling argument, this coefficient is a universal multiple of the curvature $\kappa(z^{\prime})$.
@@
This consistency means that there exists a parametrix $K(t,z,z^{\prime})$ with the specified expansions at td, sf, and ff. This parametrix solves the model problems to infinite order at td and sf and to two orders at ff. So $R(t,z,z^{\prime})=(\partial_{t}-\Delta_{\Omega})K(t,z,z^{\prime})$ has leading order 2-2, as opposed to 4-4, at ff, and leading order $\infty$ at td and sf. Using the composition formula of [10], we conclude that for each $n\geq 1$, $K\star(R^{\star n})$ has leading order at worst $t^{0}$ at ff.
@@
has order at worst 0 at ff, and since the same is true for $H^{D}_{\Omega^{\prime}}-K$ by construction of $K$, so $H^{D}_{\Omega}-H^{D}_{\Omega^{\prime}}=O(T^{0})$ at ff. The Lemma is proven. ∎
@@
As in [10], the heat trace is obtained by restricting $H^{D}_{\Omega}$ to the spatial diagonal and then integrating over $z\in\Omega$. The technical tool used is Melrose’s pushforward theorem. We will not repeat all the details here, but the upshot is that each face td, sf, and ff gives a separate contribution to the expansion (1). A priori, the expansions can interact to give logarithmic terms, but for the same reasons as in [10, p. 49], keeping in mind that we now have two orders at ff rather than one, there are no logarithmic terms until at least $O(T^{2}\log T)=O(t\log t)$ (and potentially much later, if at all). The contribution from td is
@@
Everything in these proofs works in the Neumann case as well.